OpenStudy (happpily):

simplify the number using the imaginary unit i squareroot-16

OpenStudy (happpily):

Simplify the number using the imaginary unit i \[\sqrt{-16}\]

jimthompson5910 (jim_thompson5910):

hint: \[\Large \sqrt{-16} = \sqrt{-1*16} = \sqrt{-1}*\sqrt{16}\]

jimthompson5910 (jim_thompson5910):

\[\Large i = \sqrt{-1}\]

OpenStudy (happpily):

wait isn't it 4i ?

OpenStudy (happpily):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (happpily):

alrighty thank you!

jimthompson5910 (jim_thompson5910):

np